# Continuous Product Graph Neural Networks

### v1

@@ 4.1 Experimental Stability Analysis
We generate a $600$-node product graph consisting of two connected Erdős-Rényi (ER) factor graphs with $N_{1}=20$, $N_{2}=30$, and edge probabilities $p^{(1)}_{\text{ER}}=p^{(2)}_{\text{ER}}=0.1$ for node regression. We utilize $85\%$ of the nodes in the product graph as the test set, $85\%$ of the remaining nodes as validation, and the rest of the nodes for training. We use a three-layer generative model to generate the outputs based on (6). The continuous parameters for the factor graphs are set as $t^{(1)}=2$ and $t^{(2)}=3$. The initial product graph signal $\mathbf{X}^{(0)}\in\mathbb{R}^{N\times F_{0}}$ is generated from the normal distribution with $F_{0}=6$. Each layer’s MLP of the generation model has only one layer with the number of hidden units $\{5,4,2\}$. We add noise to the factor adjacency matrices with Signal-to-Noise Ratios (SNRs) (in terms of db) of $[\infty,20,10,0,-10]$. Therefore, we construct a two-layer CITRUS and use one-layer MLP for each with $F_{1}=F_{2}=4$ to learn the representation in different SNRs to study the effect of each factor graph’s noise on the stability performance of the network.
Figure 2 shows the averaged prediction error in MSE between the true and predicted outputs over $10$ random realizations. Firstly, we notice that increasing the SNR for each factor graph improves performance, illustrating the stability properties of CITRUS. Therefore, the effect of each factor graph’s stability is well depicted in this figure, confirming the theoretical findings in Theorem 3.7, as the performance. For instance, in SNR=10 for the first graph, the performance is still severely affected by the stability for the second graph, shown by color in Figure 2. Finally, we observe that the standard deviation of the predictions is smaller for larger values of SNR, illustrating CITRUS’s robustness.
@@ 4.2 Experimental Over-smoothing Analysis
Figure 3: Oversmoothing analysis using Theorem 3.9. left: $\ln{s}-\frac{2}{P}\tilde{t}\tilde{\lambda}<0$, right: $\ln{s}-\frac{2}{P}\tilde{t}\tilde{\lambda}>0$.
For experimentally validating Theorem 3.9, we first generate a initial graph signal $\mathbf{X}_{0}\in\mathbb{R}^{N\times F_{0}}$ with $F_{0}=12$ on a $150$-node Cartesian product graph with $N_{1}=10$, $N_{2}=15$. Therefore, we consider a $10$-layer generation process based on (6) with only one-layer MLP for each layer and $\{F_{l}=12-l\}_{l=1}^{10}$, where the weight matrices are generated from normal distribution. We select ReLU as the activation function and $\{t_{l}=1\}_{l=1}^{10}$. The over-smoothing phenomenon strongly depends on the smallest spectral gap of the factor graphs as stated in Theorem 3.9. We consider two scenarios of $\ln{s}-\frac{2}{P}\tilde{t}\tilde{\lambda}<0$ and $\ln{s}-\frac{2}{P}\tilde{t}\tilde{\lambda}>0$. The first scenario relates to a well-connected product graph obtained from two factor ER graphs with $p^{(1)}_{\text{ER}}=0.05$ and $p^{(2)}_{\text{ER}}=0.95$ and, while in the other, $p^{(1)}_{\text{ER}}=p^{(2)}_{\text{ER}}=0.1$. So, we scale the MLP weight matrices by $100$ and $2.5$ to limit their maximum singular values. These settings result in $\lambda=0.03$ and $s=0.005\pm 10^{-5}$ for the first scenario; and $\lambda=0.06$ and $s=7.63\pm 10^{-6}$ for the other.
@@ 4.3 Experiments on Real-world Data
We evaluate and compare CITRUS on real-world traffic and weather spatiotemporal forecasting tasks, where we follow the settings in [10] for pre-processing and train-validation-test data partitions. All hyperparameters were optimized on the validation set. The used optimizer was Adam.
Traffic forecasting. Here, we work on two well-known datasets for spatiotemporal forecasting: MetrLA [34] and PemsBay [40]. MetrLA consists of recorded traffic data for four months on $207$ highways in Los Angeles County with $5$-minute resolutions [34]. Pems-Bay contains $5$-minute traffic load data for six months associated with $325$ stations in the Bay Area by the same setting with [40]. The spatial and temporal graphs are created by applying Gaussian kernels on the node distances [41] and simple path graphs, respectively. The task for these datasets is: given the last $30$ minutes ($T=6$ steps) of traffic recordings, we need to predict the following $15$-$30$-$60$ minutes, i.e., $H=3$, $H=6$, and $H=12$ steps ahead horizons. Our model first uses a linear encoder network, followed by three CITRUS blocks with $3$-layer MLPs for each, where $\{F^{\text{MLP}}_{l}=64\}_{l=1}^{3}$. Therefore, the output node embeddings are concatenated with the initial spatiotemporal data. Finally, we use a linear decoding layer to transform the learned representations to the number of needed horizons. We exploit residual connections within each CITRUS block to stabilize the training process [16]. We use the Mean Absolute Error (MAE) as the loss function and consider a maximum of $300$ epochs. The best model on the validation set is applied to the unseen test set. We compare CITRUS against previous methods using MAE, Mean Absolute Percentage Error (MAPE), and Root Mean Squared Error (RMSE).
@@
Weather forecasting. We also test CITRUS for weather forecasting in the Molene [42] and NOAA [43] datasets. The Molene dataset provides recorded temperature measurements for $744$ hourly intervals over $32$ measurement stations in a region in France. For the NOAA, which is associated with regions in the U.S., the temperature measurements were recorded for $8,579$ hourly intervals over $109$ stations. The pre-processing and data curation settings were similar to prior works on these datasets [10, 23]. The task here is, given the last $10$ hours of measurements, we should predict the next $\{1:5\}$ hours. For the Molene dataset, the embedding dimension and the initial linear layer of the CITRUS blocks are $16$, while no MLP modules have been used. Regarding NOAA, we set the dimension and the initial linear layer of the CITRUS blocks as $16$, and use $3$-layer MLPs for channel mixing in each block. The CITRUS blocks have the last activation as a Leaky ReLU. We use MSE as the loss function and consider root-normalized MSE (rNMSE) as the evaluation metric.
@@ 5 Conclusion and Limitations
An interesting future direction for our framework could involve adapting it to handle other types of graph products, such as Kronecker and strong graph products [22]. Future efforts will also focus on finding tighter and more general upper bounds in the stability and over-smoothing analyses. For example, we will explore the possible relationships between the size of the factor graphs and these properties, either in a general form or for specific well-studied structures, such as ER graphs. Additionally, we plan to investigate more challenging real-world applications of the proposed framework, especially in scenarios involving more than two factor graphs.
6 Acknowledgement
@@ Appendix A Proofs
Now, by performing mode-$r$ unfolding operation on $\underline{\mathbf{U}}_{t,p}$, we have:
@@
Using the pairwise summation property of the factor eigenvalues in a Cartesian product, one can write
@@
By construction, we first prove the theorem for the case of $P=2$. The normalized Laplacian $\hat{\mathbf{L}}$ can be rewritten as:
@@
$E(\mathbf{x})=\mathbf{x}^{\top}\hat{\mathbf{L}}\mathbf{x}=\sum_{i=1}^{N}{\tilde{x}_{i}\lambda^{2}_{i}}.$ (33)
Next, by defining $\lambda$ as the smallest non-zero eigenvalue of the Laplacian $\hat{\mathbf{L}}$, the following lemma characterizes the over-smoothing aspects of applying a heat kernel in the simplest case:
@@
Let $s\vcentcolon=\sup_{l\in\mathbb{N}_{+}}{s_{l}}$. We have
@@
Appendix B Descriptions of the compared methods
@@
Appendix C Experiments on more than two-factor graphs
We consider the node regression task similar to the settings described in Section 4.1 but for three ER graphs with $p^{(1)}_{\text{ER}}=p^{(1)}_{\text{ER}}=0.3$, and varying $p^{(3)}_{\text{ER}}\in\{0.1,0.3\}$ to monitor the performance in two different connectivity scenarios. Note that here we do not perturb factor graphs but, to make the experiment more challenging, the SNR on the graph data is set to SNR=0. Table 5 provides the results across a varying number of layers, which shows that the proposed framework’s performance is resistant to adding layers, especially in the case of more sparse graphs, i.e., $p^{(3)}_{\text{ER}}=0.1$. On the other hand, the best results were obtained in $l=1,2,4$, which are the nearest numbers to the actual number of layers in the true generative process which was 3. These observations are validated by comparing the results with the GCN, in which the over-smoothing and performance degradation happen severely faster than CITRUS.
Table 5: Experiments on more than two factor graphs ($p^{(1)}\_{\text{ER}}=p^{(2)}\_{\text{ER}}=0.3$).
$l=1$ $l=2$ $l=4$ $l=8$ $l=16$ $l=32$
GCN, $p^{(3)}\_{\text{ER}}=0.1$ 0.185$\pm$0.132 0.241$\pm$0.096 0.371$\pm$0.132 0.458$\pm$0.132 0.494$\pm$0.147 0.552$\pm$0.118
CITRUS, $p^{(3)}\_{\text{ER}}=0.1$ 0.141$\pm$0.075 0.137$\pm$0.082 0.150$\pm$0.091 0.198$\pm$0.106 0.215$\pm$0.124 0.276$\pm$0.120
GCN, $p^{(3)}\_{\text{ER}}=0.3$ 0.253$\pm$0.211 0.296$\pm$0.113 0.365$\pm$0.123 0.429$\pm$0.097 0.517$\pm$0.143 0.67$\pm$0.153
CITRUS, $p^{(3)}\_{\text{ER}}=0.3$ 0.188$\pm$0.094 0.182$\pm$0.095 0.193$\pm$0.096 0.230$\pm$0.089 0.242$\pm$0.096 0.336$\pm$0.107
@@ Appendix D Standard deviation of the results
Appendix E Experiments Compute Resources

### v2

@@ 4.1 Experimental Stability Analysis
We generate a $600$-node product graph consisting of two connected Erdős-Rényi (ER) factor graphs with $N_{1}=20$, $N_{2}=30$, and edge probabilities $p^{(1)}_{\text{\tiny ER}}=p^{(2)}_{\text{\tiny ER}}=0.1$ for node regression. We utilize $15\%$ of the nodes in the product graph as the test set, $15\%$ of the remaining nodes as validation, and the rest of the nodes for training. We use a three-layer generative model to generate the outputs based on (6). The continuous parameters for the factor graphs are set as $t^{(1)}=2$ and $t^{(2)}=3$. The initial product graph signal $\mathbf{X}^{(0)}\in\mathbb{R}^{N\times F_{0}}$ is generated from the normal distribution with $F_{0}=6$. Each layer’s MLP of the generation model has only one layer with the number of hidden units $\{5,4,2\}$. We add noise to the factor adjacency matrices with signal-to-noise ratios (SNRs) (in terms of db) of $\{\infty,20,10,0,-10\}$. Therefore, we construct a two-layer CITRUS and use one-layer MLP for each with $F_{1}=F_{2}=4$ to learn the representation in different SNRs to study the effect of each factor graph’s noise on the stability performance of the network.
Figure 2 shows the averaged prediction error in MSE between the true and predicted outputs over $10$ random realizations. Firstly, we notice that increasing the SNR for each factor graph improves performance, illustrating the stability properties of CITRUS. Therefore, the effect of each factor graph’s stability on the overall stability is well depicted in this figure, confirming the theoretical findings in Theorem 3.7. For instance, in SNR=10 for the first graph, the performance is still severely affected by the stability for the second graph, shown by color in Figure 2. Finally, we observe that the standard deviation of the predictions is smaller for larger values of SNR, illustrating CITRUS’s robustness.
@@ 4.2 Experimental Over-smoothing Analysis
Figure 3: Over-smoothing analysis using Theorem 3.10. left: $\ln{s}-\frac{2}{P}\tilde{t}\tilde{\lambda}<0$, right: $\ln{s}-\frac{2}{P}\tilde{t}\tilde{\lambda}>0$.
For experimentally validating Theorem 3.10, we first generate a initial graph signal $\mathbf{X}_{0}\in\mathbb{R}^{N\times F_{0}}$ with $F_{0}=12$ on a $150$-node Cartesian product graph with $N_{1}=10$, $N_{2}=15$. Therefore, we consider a $10$-layer generation process based on (6) with only one-layer MLP for each layer and $\{F_{l}=12-l\}_{l=1}^{10}$, where the weight matrices are generated from normal distribution. We select ReLU as the activation function and $\{t_{l}=1\}_{l=1}^{10}$. The over-smoothing phenomenon strongly depends on the smallest spectral gap of the factor graphs as stated in Theorem 3.10. We consider two scenarios of $\ln{s}-\frac{2}{P}\tilde{t}\tilde{\lambda}<0$ and $\ln{s}-\frac{2}{P}\tilde{t}\tilde{\lambda}>0$. The first scenario relates to a well-connected product graph obtained from two factor ER graphs with $p^{(1)}_{\text{\tiny ER}}=0.05$ and $p^{(2)}_{\text{\tiny ER}}=0.95$ and, while in the other, $p^{(1)}_{\text{\tiny ER}}=p^{(2)}_{\text{\tiny ER}}=0.1$. So, we scale the MLP weight matrices by $100$ and $2.5$ to limit their maximum singular values. These settings result in $\lambda=0.06$ and $s=0.004\pm 10^{-5}$ for the first scenario; and $\lambda=0.07$ and $s=7.691\pm 10^{-6}$ for the other.
@@ 4.3 Experiments on Real-world Data
We evaluate and compare CITRUS on real-world traffic and weather spatiotemporal forecasting tasks, where we follow the settings in [10] for pre-processing and training-validation-testing data partitions. All hyperparameters were optimized on the validation set with the details in Section I. We use Adam as optimizer.
Traffic forecasting. Here, we work on two well-known datasets for spatiotemporal forecasting: MetrLA [35] and PemsBay [36]. MetrLA consists of recorded traffic data for four months on $207$ highways in Los Angeles County with $5$-minute resolutions [35]. PemsBay contains $5$-minute traffic load data for six months associated with $325$ stations in the Bay Area by the same setting with [36]. The spatial and temporal graphs are created by applying Gaussian kernels on the node distances [37] and simple path graphs, respectively. The task for these datasets is, by having the last $30$ minutes ($T=6$ steps) of traffic recordings, we need to predict the following $15$-$30$-$60$ minutes, i.e., $H=3$, $H=6$, and $H=12$ steps ahead horizons. Our model first uses a linear encoder network, followed by three CITRUS blocks with $3$-layer MLPs for each, where $\{F^{\text{MLP}}_{l}=64\}_{l=1}^{3}$. Therefore, the output node embeddings are concatenated with the initial spatiotemporal data. Finally, we use a linear decoding layer to transform the learned representations to the number of needed horizons. We exploit residual connections within each CITRUS block to stabilize the training process [17]. We use the mean absolute error (MAE) as the loss function and consider a maximum of $300$ epochs. The best model on the validation set is applied to the unseen test set. We compare CITRUS against previous methods using MAE, mean absolute percentage error (MAPE), and root mean squared error (RMSE).
@@
Weather forecasting. We also test CITRUS for weather forecasting in the Molene [43] and NOAA [44] datasets. The Molene dataset provides recorded temperature measurements for $744$ hourly intervals over $32$ measurement stations in a region in France. For the NOAA, which is associated with regions in the U.S., the temperature measurements were recorded for $8,579$ hourly intervals over $109$ stations. The pre-processing and data curation settings were similar to prior works on these datasets [10, 25]. The task here is by having the last $10$ hours of measurements, we should predict the next $\{1:5\}$ hours. For the Molene dataset, the embedding dimension and the initial linear layer of the CITRUS blocks are $16$, while no MLP modules have been used. Regarding NOAA, we set the dimension and the initial linear layer of the CITRUS blocks as $16$, and use $3$-layer MLPs for channel mixing in each block. The CITRUS blocks have the last activation as a Leaky ReLU. We use MSE as the loss function and consider root-normalized MSE (rNMSE) as the evaluation metric.
@@ 5 Conclusion and Limitations
An interesting future direction for our framework could involve adapting it to handle other types of graph products, such as Kronecker and Strong graph products [23, 24]. Future efforts will also focus on finding tighter and more general upper bounds in the stability and over-smoothing analyses. For example, we will explore the possible relationships between the size of the factor graphs and these properties, either in a general form or for specific well-studied structures, such as ER graphs. Additionally, we plan to investigate more challenging real-world applications of the proposed framework, especially in scenarios involving more than two factor graphs.
Acknowledgments
@@
Appendix
The appendix contains the proof of the claimed theoretical statements (Section A), descriptions of the compared methods (Section C), the standard deviation of the results (Section D), homophily-heterophily trade-off (Section E), the case of inexact Cartesian product graphs (Section F), how to choose the appropriate number of eigenvalue-eigenvector pairs (Section G), the effect of graph receptive fields on the over-smoothing (Section H) and hyperparameter details (Section I), respectively.
@@ Appendix A Proofs
Now, by performing mode-$p$ unfolding operation on $\underline{\mathbf{U}}_{t,p}$, we have:
@@
Using the pairwise summation property of the factor eigenvalues in a Cartesian product, one can write:
@@
By construction, we first prove the theorem for the case of $P=2$. The defined Laplacian $\hat{\mathbf{L}}$ in (15) can be rewritten as:
@@
$E(\mathbf{x})=\mathbf{x}^{\top}\hat{\mathbf{L}}\mathbf{x}=\sum_{i=1}^{N}{\lambda_{i}\tilde{x}^{2}_{i}}.$ (33)
Next, by defining $\lambda$ as the smallest non-zero eigenvalue of the Laplacian $\hat{\mathbf{L}}$, the following lemma characterizes the over-smoothing aspects of applying a heat kernel in the simplest case.
@@
Let $s\vcentcolon=\sup_{l\in\mathbb{N}_{+}}{s_{l}}$. We have:
@@
Appendix B Experiments on more than Two Factor Graphs
We consider the node regression task similar to the settings described in Section 4.1 but for three ER graphs with $p^{(1)}_{\text{\tiny ER}}=p^{(1)}_{\text{\tiny ER}}=0.3$, and varying $p^{(3)}_{\text{\tiny ER}}\in\{0.1,0.3\}$ to monitor the performance in two different connectivity scenarios. Note that here we do not perturb factor graphs but, to make the experiment more challenging, the SNR on the graph data is set to SNR=0. Table 5 provides the results across a varying number of layers, which shows that the proposed framework’s performance is resistant to adding layers, especially in the case of more sparse graphs, i.e., $p^{(3)}_{\text{\tiny ER}}=0.1$. On the other hand, the best results were obtained in $l=1,2,4$, which are the nearest numbers to the actual number of layers in the true generative process which was 3. These observations are validated by comparing the results with the GCN, in which the over-smoothing and performance degradation happen severely faster than CITRUS.
Table 5: Experiments on more than two factor graphs ($p^{(1)}_{\text{\tiny ER}}=p^{(2)}_{\text{\tiny ER}}=0.3$).
$l=1$ $l=2$ $l=4$ $l=8$ $l=16$ $l=32$
GCN, $p^{(3)}_{\text{\tiny ER}}=0.1$ 0.185$\pm$0.132 0.241$\pm$0.096 0.371$\pm$0.132 0.458$\pm$0.132 0.494$\pm$0.147 0.552$\pm$0.118
CITRUS, $p^{(3)}_{\text{\tiny ER}}=0.1$ 0.141$\pm$0.075 0.137$\pm$0.082 0.150$\pm$0.091 0.198$\pm$0.106 0.215$\pm$0.124 0.276$\pm$0.120
GCN, $p^{(3)}_{\text{\tiny ER}}=0.3$ 0.253$\pm$0.211 0.296$\pm$0.113 0.365$\pm$0.123 0.429$\pm$0.097 0.517$\pm$0.143 0.67$\pm$0.153
CITRUS, $p^{(3)}_{\text{\tiny ER}}=0.3$ 0.188$\pm$0.094 0.182$\pm$0.095 0.193$\pm$0.096 0.230$\pm$0.089 0.242$\pm$0.096 0.336$\pm$0.107
Appendix C Descriptions of the Compared Methods
@@ Appendix D Standard Deviation of the Results
Appendix E Homophily-Heterophily Trade-off across the Studied Real Datasets
Regarding the homophily-heterophily trade-off, first, we should mention that measuring the homophily-heterophily indices is way more challenging in spatiotemporal settings rather than in single-graph scenarios [46], since we are facing two intra-graph (i.e., the spatial dimension) and inter-graph (i.e., the temporal domain of changing time series characteristics) aspects. In this way, in the recent pioneer literature, they considered MetrLA, PemsBay, and temperature datasets (like Molene and NOAA used in our paper) as complicated in terms of homophily-heterophily behaviors. We provided the measured homophily-heterophily metrics (full details in [46]) in Table 8. As observed in this table, due to the wide range of variability across the used datasets, we have actually evaluated our performance on various datasets in terms of homophily-heterophily levels. For instance, Molene acts severely heterophilic in both intra and inter-graph domains, while NOAA acts more homophilic than Molene in the inter-graph domain.
Table 8: Intra and Inter-homophily measures on the real-world graphs of MetrLA and PemsBay datasets.
Dataset Intra-graph homophily $\boldsymbol{\pi}^{s}_{v_{i}}$ Inter-graph homophily $\boldsymbol{\pi}^{T}_{v_{i}}$
$p_{s}$ $q_{p}$ $q_{n}$ $p_{s}$ $q_{p}$ $q_{n}$
MetrLA 0.2273 0.4732 0.2995 0.3325 0.4920 0.1755
PemsBay 0.1073 0.5912 0.3015 0.2399 0.6863 0.0738
Molene 0.0148 0.6248 0.3602 0.0152 0.5545 0.4301
NOAA 0.1249 0.5345 0.3405 0.2862 0.4351 0.2786
Appendix F The Case of Inexact Cartesian Product Modeling
This scenario lies under the general umbrella of inaccurate adjacencies [30], which has been rigorously studied in the stability analysis in Section 3.2. However, in a simplified and insightful case of facing graph products other than Cartesian, we consider two well-known graph products, i.e., Strong and Kronecker products [24, 23]. Consider the true graph is actually a Strong product graph $\mathbf{A}=\mathbf{A}_{1}\otimes\mathbf{A}_{2}+\mathbf{A}_{1}\otimes\mathbf{I}_{2}+\mathbf{I}_{1}\otimes\mathbf{A}_{2}$, and one mistakenly considers it as Cartesian product $\tilde{\mathbf{A}}=\mathbf{A}_{1}\otimes\mathbf{I}_{2}+\mathbf{I}_{1}\otimes\mathbf{A}_{2}$. Then, $\mathbf{E}=\mathbf{A}-\tilde{\mathbf{A}}=\mathbf{A}_{1}\otimes\mathbf{A}_{2}$, where $\|\mathbf{E}\|=\lambda^{(1)}_{max}\lambda^{(2)}_{max}$. So, the approximation error depends on the multiplication of the factor graph eigenspaces. Similarly, in the case of Kronecker product graphs $\tilde{\mathbf{A}}=\mathbf{A}_{1}\otimes\mathbf{A}_{2}$, it obtains $\|\mathbf{E}\|\leq\lambda^{(1)}_{max}+\lambda^{(2)}_{max}+\lambda^{(1)}_{max}\lambda^{(2)}_{max}$. So, in this case, the summation of the factor eigenspaces matters too. Briefly, the approximation error bound depends on the factor graph spectrums.
Appendix G Choosing Appropriate $k$
We can choose an appropriate $k$ using two methodologies: supervised and unsupervised. For the supervised option, we can use cross-validation, as we did in our results. For the unsupervised method, we can analyze the Laplacian matrices for a low-rank approximation task. For example, in Figure 4, we plot the explained variances of the principal components of the spatial Laplacian (i.e., $\|\mathbf{v}\lambda\mathbf{v}^{\top}\|_{F}^{2}$ for $\mathbf{v}$ and $\lambda$ being the eigenvectors and eigenvalues) on the MetrLA dataset. We observe a strong concentration of variance in a few eigenvalues. We also observe that we can capture almost 80% of the explained variance by only relying on 50 components.
Figure 4: Explained variance ratio vs. selected principal components.
Appendix H The Effect of the Graph Receptive Fields on Controlling Over-smoothing
From a theoretical point of view, based on eq. (18), one intuitive way to alleviate or at least slow down the over-smoothing phenomena is to keep $\ln{s}-\frac{2}{P}\tilde{t}\tilde{\lambda}$ positive, which tends to $\tilde{t}<\frac{P}{2\tilde{\lambda}}\ln{s}$. This clarifies why one should not increase the node-neighborhood especially in the case of facing a connected graph or increasing the number of layers. This finding is also compatible with the pioneering research about the over-smoothing concept, e.g., [14]. From an experimental point of view, and if we consider the graph receptive field $t$ as a hyperparameter, we have designed an experiment to validate the claimed statement. Here, we vary $t\in\{0.1,1,5,10,20\}$ and monitor the over-smoothing process in Figure 5. In this figure, $b=\frac{P}{2\tilde{\lambda}}\ln{s}$, and as it is observed, for values of $t$ higher than $b$ the over-smoothing phenomena with increasing the number of layers is happening way faster than the other ones, which validates the theoretical discussion, too. Apart from this, the weight normalization technique [14, 15] has also been mentioned as a helpful action but it hinders the theoretical foundations of CITRUS. Note that, in practice, only escaping from the over-smoothing is not enough to have good performance. Therefore, for $t<b$, one might need to perform other kinds of analyses, e.g., over-squashing analysis.
Appendix I Hyperparameter Details
The detailed hyperparameters (optimized by cross-validation on the validation data) and/or training settings of CITRUS , i.e., $T$ (auto-regressive order), $emb$ (dimension of embedding size in the spatiotemporal encoder), $hid$ (dimension of linear mapping size in the spatiotemporal encoder), $F_{\text{MLP}}$ (dimension of linear mapping size in the MLP layers), $n_{\text{CITRUS}}$ (number of CITRUS blocks), $F$ (second dimension of $\mathbf{W}_{l}$ in (6)), $k_{i}$ (number of selected eigenvalue-eigenvector pairs in the $i$-th factor graph), $lr$ (learning rate), $n_{\text{batch}}$ (batch size), and $n_{\text{epochs}}$ (number of epochs), are provided in Table 9. Full details can be found in the implementation codes in https://github.com/ArefEinizade2/CITRUS.
Figure 5: Log relative distance vs. increasing the number of layers for different values of $t$ in actual bounds.
Table 9: Details of the training settings and hyperparameters, i.e., $T$ (auto-regressive order), $emb$ (dimension of embedding size in the spatiotemporal encoder), $hid$ (dimension of linear mapping size in the spatiotemporal encoder), $F_{\text{MLP}}$ (dimension of linear mapping size in the MLP layers), $n_{\text{CITRUS}}$ (number of CITRUS blocks), $F$ (second dimension of $\mathbf{W}_{l}$ in (6)), $k_{i}$ (number of selected eigenvalue-eigenvector pairs in the $i$-th factor graph), $lr$ (learning rate), $n_{\text{batch}}$ (batch size), and $n_{\text{epochs}}$ (number of epochs).
Dataset $T$ $emb$ $hid$ $F_{\text{MLP}}$ $n_{\text{CITRUS}}$ $F$ $k_{1}$ $k_{2}$ $lr$ $n_{\text{batch}}$ $n_{\text{epochs}}$
MetrLA 6 16 32 64 3 64 205 4 0.01 2048 300
PemsBay 6 16 32 64 3 64 323 4 0.01 2048 300
Molene 10 16 16 16 3 16 30 8 0.001 64 1500
NOAA 10 4 4 4 3 4 107 8 0.01 256 400
Appendix J Experiments Compute Resources